Round to four decimal places as needed.) ol
Answer #1

a) P(8 men and 1 women)

50) /10 60

= 0.3632

b) P(7 men and 2 women)

= \frac{\binom{50}{7}\binom{10}{2}}{\binom{60}{9}}

= 0.3040

c) Option D is correct.

d) P(Atleast one women)

= 1 - P(All men)

= 1-\frac{\binom{50}{9}}{\binom{60}{9}}

= 0.8305